OpenStudy (anonymous):

standard form of the line which passes through (1,-3) ad as a slope of 2

hartnn (hartnn):

lets take slope point form, the we'll convert to standard form the slope point form is (y-y1) = m (x-x1) where m=2 is slope and (x1,y1) is the point on line so can u tell me equation of line in slope point form first ?

hartnn (hartnn):

?

OpenStudy (anonymous):

would it be 2x-y=-5?

hartnn (hartnn):

y+3=2(x-1) y+3=2x-2 2x-y=3+2=5

OpenStudy (anonymous):

like the anwser be what I just said ?

hartnn (hartnn):

2x-y=5 is the answer

OpenStudy (anonymous):

thank you